# Liquid Resistance Liquid Capacitance Networks

### v2

@@ A.1.1 Proof of the Smoothness Theorem
A.1.2 Proof of the Accuracy Theorem
Theorem 2 (LRC Accuracy).
Let $h^{s}$ be an STC-model instance of Equation (2) with $h^{r}$ being an associated LRC-model instance of Equation (3) with $\lambda^{r}<\lambda^{s}$ (Theorem. 1). Let $o^{T}_{t}$ be training labels and $\hat{o}^{T,r}_{t}$ and $\hat{o}^{T,s}_{t}$ be the predictions of the models $h^{r}$ and $h^{s}$ respectively. Let both models have a small training loss $\|o^{T}_{t}-\hat{o}^{T}_{t}\|_{2}<\epsilon_{T}$ with small $\epsilon_{T}>0$. Let $y^{T}_{t}$ and $y^{V}_{t}$ be respectively training and validation input. Let the validation set be similar to the training set in a sense that for every time $t$ it holds that $\|y^{T}_{t}-y^{V}_{t}\|_{2}<\epsilon_{y}$ and $\|o^{T}_{t}-o^{V}_{t}\|_{2}<\epsilon_{o}$ with small $\epsilon_{y},\epsilon_{o}>0$. Then it holds that the upper bound of the validation loss of the LRC-model is smaller than the one of the STC-model.
@@ A.1.2 Proof of the Accuracy Theorem
$\leq\frac{1}{T}\sum_{t=1}^{T}\epsilon_{o}+\epsilon_{T}+\|\hat{o}^{T}_{t}-\hat{o}^{V}_{t}\|_{2}$ (44)
@@
$\leq\epsilon_{y}\cdot\lambda,$ (49)
@@
$L(o^{V},\hat{o}^{V})\leq\epsilon_{o}+\epsilon_{T}+\epsilon_{y}\cdot\lambda=\bar{L}(o^{V},\hat{o}^{V})$ (50)
Finally, by using Theorem 1, it holds that:
$\bar{L}(o^{V},\hat{o}^{V,r})=\epsilon_{o}+\epsilon_{T}+\epsilon_{y}\cdot\lambda^{r}<\epsilon_{o}+\epsilon_{T}+\epsilon_{y}\cdot\lambda^{s}=\bar{L}(o^{V},\hat{o}^{V,s})$ (51)
This proves that the validation loss of LRCs is smaller then the validation loss of STCs. ∎
A.2 Liquid Resistance Liquid Capacitance Units (LRCU)
In this section we first introduce the RNN formulation of LRCUs, which as discussed in the paper, are a very accurate and efficient Euler integration of order one of LRCs, with a time step of one. Since LRCUs turn out to be a new form of gated recurrent units, we explore their relation to GRUs.
A.2.1 LRCUs as an RNN
@@
$\begin{array}[]{c}h_{i,t}=(1-\epsilon(w_{i,t})\,\sigma(f_{i,t}))\,h_{i,t-1}+\epsilon(w_{i,t})\,\tau(u_{i,t})\,e_{li}\\[5.69054pt] f_{i,t}=\sum_{j=1}^{m+n}g_{ji}\sigma(a_{ji}y_{j,t}+b_{ji})+g_{li}\\[5.69054pt] u_{i,t}=\sum_{j=1}^{m+n}k_{ji}\sigma(a_{ji}y_{j,t}+b_{ji})+g_{li}\\[5.69054pt] w_{i,t}=\sum_{j=1}^{m+n}o_{ji}y_{j,t}+p_{i}\end{array}$ (52)
@@
A.2.2 LRCUs versus GRUs
In this section we explore the connections between LRCUs and GRUs. The general form of a gated recurrent units (GRU) is according to Cho et al. (2014), an RNN of the following form:
$\begin{array}[]{c}h_{i,t}=(1-\sigma(f_{i,t}))\,h_{i,t-1}+\sigma(f_{i,t})\,\tau(u_{i,t})\\[8.53581pt] f_{i,t}=\sum_{j=1}^{m+n}a^{f}_{ji}\,y_{j,t}+b^{f}_{j}\\[5.69054pt] u_{i,t}=\sum_{j=1}^{m+n}a^{u}_{ji}\,y^{\prime}_{j,t}+b^{u}_{j}\\[5.69054pt] r_{i,t}=\sum_{j=1}^{m+n}a^{r}_{ji}\,y_{j,t}+b^{r}_{j}\end{array}$ (53)
@@ A.2.2 LRCUs versus GRUs
$\begin{array}[]{c}\dot{h}_{i}=\sigma(f_{i})(-h_{i}+\tau(u_{i}))\\[5.69054pt] f_{i}=\sum_{j=1}^{m+n}a^{f}_{ji}\,y_{j}+b^{f}_{j}\\[5.69054pt] u_{i}=\sum_{j=1}^{m+n}a^{u}_{ji}\,y^{\prime}_{j}+b^{u}_{j}\\[5.69054pt] r_{i}=\sum_{j=1}^{m+n}a^{r}_{ji}\,y_{j}+b^{r}_{j}\end{array}$ (54)
Here the vector $y\,{=}\,[h,x]$ is defined as before, and vector $y^{\prime}\,{=}\,[\sigma(r)\odot{}h,x]$. The RG occurring in $y^{\prime}$ determines how state $h$ is used in the update part $\tau(u_{i})$ of the ODE. The LRC’s time constant (TC) thus consists of a liquid resistance liquid capacitance, while the GRU’s TC consists of a liquid capacitance, only. As a consequence, GRUs have a less expressive TC but more expressive update.
A.3 Experiments for LRCs and LRCUs
In this section, we provide additional implementation details about our experimental evaluation of LRC in ODE modeling and LRCUs on the popular benchmarks for gated recurrent units. The benchmarks considered were the Localization Data for Person Activity, the IMDB Movie Sentiment Classification, the Permuted Sequential MNIST, and a relatively complex autonoumous-driving in the Lane-Keeping Task.
In general, LRC(u)s have more trainable parameters per cell. Therefore, we are training the other models with more cells to ensure a fair comparison between them.
The experiments are run on an Nvidia Tesla T4.
A.3.1 Neural ODE Experiments
For both the linear and non-linear ODE modeling task, we used a 3-layer (units of 32, 32 and 2) Neural ODE and the LRC (units of 16, 16, 2), which hold 1k trainable parameters. The hyperparameters are presented in Table 3. We use sequences of 16 points during training and test the models on the whole sequences of 1000 data points giving them only the initial state information of $(x_{0},y_{0})$. As the non-linear ODE modeling task is more challenging, we allocated double the training time compared to the linear spiral ODE task. The training time is 0.4-0.5 seconds per iteration for both models.
@@ A.3.1 Neural ODE Experiments
Iterations $1000/2000$
A.3.2 Localization Data for Person Activity
@@
For LSTMs, GRUs, and MGUs we concatenate the time-step information (difference in time between to inputs) directly with the input features. This time step is used as the $\Delta_{t}$ value in LRCUs. LSTM, GRU and MGU networks contain 100 units, while LRCUs have 100 units.
@@
A.3.3 IMDB Movie Sentiment Classification
@@
Parameters 40k 30k 20k 20k 20k
A.3.4 Permuted Sequential MNIST
@@
During the training of this task, the computed loss values of MGUs became NaNs in 2 out of the 3 runs in the middle of the experiments. Thus, for plotting the validation accuracy in Figure 5, we included only the one fully successful run for them.
@@
A.3.5 Lane-Keeping Task
@@
Table 10: The shape and size of the layers in the convolutional head of the RNN-policy used to solve the Lane-Keeping Task. Settings are adapted from Farsang et al. (2023).
@@
Conv2D Filters: 24, Kernel size: 5, Stride: 2, Activ.: ReLU
Conv2D Filters: 36, Kernel size: 5, Stride: 1, Activ.: ReLU
@@
Conv2D Filters: 48, Kernel size: 3, Stride: 1, Activ.: ReLU
@@
Conv2D Filters: 64, Kernel size: 3, Stride: 1, Activ.: ReLU
@@
Conv2D Filters: 64, Kernel size: 3, Stride: 1, Activ.: ReLU
@@
Table 11: Hyperparameters of the Lane-Keeping experiment.
@@
Optimizer AdamW (Loshchilov and Hutter, 2017)
@@
A.4 Interpretability Experiments for LRCUs
In this section we provide additional details for the interpretability experiments for the Lane-Keeping Task. In Figure 12, we provide the attention maps for the winter season.
@@ A.4 Interpretability Experiments for LRCUs
In Figure 13 we provide the neural activity of two cells of the lane-keeping policy for the winter season. As one can clearly see, the LRCU-S has a very smooth behavior closely matching the road geometry: Cell 1 is responsible for turning left, and Cell 2 is responsible for turning right.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Interpretability. The neural activity of two cells in the learned policy for the Lane-Keeping Task, projected over time on the 1km road driven in winter, for three popular gated recurrent units (LSTMs, GRUs, and MGUs) and for LRC units (LRCU-A, and LRCU-S). It is very hard to visually match the neural activity of LSTM, GRU, and MGU cells to the traversed road geometry, respectively. However, LRCU-S cells especially, display a very smooth and identifiable pattern during driving: Cell 1 is responsible for turning right, and Cell 2 for turning left.
In Table 13, we show the absolute correlation between the neural activity of all cells and the trajectory of the road in winter. As one can see, in the case of LRCU-S, they are highly correlated.
Table 13: Interpretability. Absolute correlation between the neural activity and the trajectory of the road in winter. Values closer to 1 indicate strong correlation. Results are averaged over 3 runs.
LSTM GRU MGU LRCU-A LRCU-S
Winter $0.450\pm 0.262$ $0.389\pm 0.248$ $0.308\pm 0.232$ $0.533\pm 0.292$ $\mathbf{0.766\pm 0.243}$
Finally, in Figure 14 we depict the robustness of attention, measured by the structural-similarity index, in summer and winter. As one can see, the LRCU-S has the highest index among all RNNs.
Figure 14: Robustness of the attention, measured by the Structural Similarity Index (SSIM) of the models, in summer (left) and in winter (right). The LRCU models maintain the most similar focus of their attention in the presence of noise, indicated by the boxplots closer to 1. Lighter color refers to additional Gaussian noise of zero-mean and $\sigma^{2}=0.1$ variance, and darker to $\sigma^{2}=0.2$ variance.
By measuring the change of attention, we also compute the Structural Similarity Index (SSIM) (Wang et al., 2004) of the images pairwise, between the noise-free and the noisy saliency map for each model, respectively. This technique assesses the similarity between two images, with a focus on image degradation. An SSIM value of one denotes full similarity, whereas one of zero, denotes zero similarity. In our particular case, where we are comparing attention-map images, values close to one mean that the extra noise leads to less distortion in the attention map, which is a very desirable aim in the development of robust controllers. The SSIM value for $1600{-}1600$ pairs of images, used in the attention-map comparisons between noise-free and noisy, summer and winter images, respectively, are presented in Figure 14. All results in this table are averaged over three runs for each model. As one can see, LRCUs demonstrate much more robust attention in the presence of Gaussian noise.

### v4

@@ 0.A.1.1 Proof of the Lipschitz constant Theorem
0.A.1.2 Proof of the Generalization Theorem
Theorem 0.A.2 (LRC Generalization).
Let $h^{s}$ be an STC-model instance of Equation (2) with $h^{r}$ being an associated LRC-model instance of Equation (3) with $\lambda^{r}<\lambda^{s}$ (Theorem. 3.1). Let $o^{T}_{t}$ be training labels and $\hat{o}^{T,r}_{t}$ and $\hat{o}^{T,s}_{t}$ be the predictions of the models $h^{r}$ and $h^{s}$ respectively. Let both models have a small training loss $\|o^{T}_{t}-\hat{o}^{T}_{t}\|_{2}<\epsilon_{T}$ with small $\epsilon_{T}>0$. Let $y^{T}_{t}$ and $y^{V}_{t}$ be respectively training and validation input. Let the validation set be similar to the training set (drawn in-distribution) in a sense that for every time $t$ it holds that $\|y^{T}_{t}-y^{V}_{t}\|_{2}<\epsilon_{y}$ and $\|o^{T}_{t}-o^{V}_{t}\|_{2}<\epsilon_{o}$ with small $\epsilon_{y},\epsilon_{o}>0$. Then it holds that the upper bound of the validation loss of the LRC-model is smaller than the one of the STC-model.
@@ 0.A.1.2 Proof of the Generalization Theorem
$\leq\frac{1}{T}\sum_{t=1}^{T}\epsilon_{o}+\epsilon_{T}+\|\hat{o}^{T}_{t}-\hat{o}^{V}_{t}\|_{2}.$ (51)
@@
$\leq\epsilon_{y}\cdot\lambda.$ (56)
@@
$L(o^{V},\hat{o}^{V})\leq\epsilon_{o}+\epsilon_{T}+\epsilon_{y}\cdot\lambda=\bar{L}(o^{V},\hat{o}^{V}).$ (57)
Finally, by using Theorem 3.1, it holds that
$\bar{L}(o^{V},\hat{o}^{V,r})=\epsilon_{o}+\epsilon_{T}+\epsilon_{y}\cdot\lambda^{r}$ (58)
$<\epsilon_{o}+\epsilon_{T}+\epsilon_{y}\cdot\lambda^{s}=\bar{L}(o^{V},\hat{o}^{V,s}).$ (59)
This proves that the validation loss of LRCs is smaller than the validation loss of STCs. ∎
0.A.2 Liquid-Resistance Liquid-Capacitance Units (LRCU)
In this section, we first introduce the RNN formulation of LRCUs, which as discussed in the paper, are a very accurate and efficient Euler integration of order one of LRCs, with a time step of one. Since LRCUs turn out to be a new form of gated recurrent units, we explore their relation to GRUs.
0.A.2.1 LRCUs as an RNN
@@
$\begin{array}[]{c}h_{i,t}=(1-\epsilon(w_{i,t})\,\sigma(f_{i,t}))\,h_{i,t-1}+\epsilon(w_{i,t})\,\tau(u_{i,t})\,e_{li}\\[5.69054pt] f_{i,t}=\sum_{j=1}^{m+n}g_{ji}\sigma(a_{ji}y_{j,t}+b_{ji})+g_{li}\\[5.69054pt] u_{i,t}=\sum_{j=1}^{m+n}k_{ji}\sigma(a_{ji}y_{j,t}+b_{ji})+g_{li}\\[5.69054pt] w_{i,t}=\sum_{j=1}^{m+n}o_{ji}y_{j,t}+p_{i}.\end{array}$ (60)
@@
0.A.2.2 LRCUs versus GRUs
In this section we explore the connections between LRCUs and GRUs. The general form of a gated recurrent units (GRU) is according to [8], an RNN of the following form:
$\begin{array}[]{c}h_{i,t}=(1-\sigma(f_{i,t}))\,h_{i,t-1}+\sigma(f_{i,t})\,\tau(u_{i,t})\\[8.53581pt] f_{i,t}=\sum_{j=1}^{m+n}a^{f}_{ji}\,y_{j,t}+b^{f}_{j}\\[5.69054pt] u_{i,t}=\sum_{j=1}^{m+n}a^{u}_{ji}\,y^{\prime}_{j,t}+b^{u}_{j}\\[5.69054pt] r_{i,t}=\sum_{j=1}^{m+n}a^{r}_{ji}\,y_{j,t}+b^{r}_{j}.\end{array}$ (61)
@@ 0.A.2.2 LRCUs versus GRUs
$\begin{array}[]{c}\dot{h}_{i}=\sigma(f_{i})(-h_{i}+\tau(u_{i}))\\[5.69054pt] f_{i}=\sum_{j=1}^{m+n}a^{f}_{ji}\,y_{j}+b^{f}_{j}\\[5.69054pt] u_{i}=\sum_{j=1}^{m+n}a^{u}_{ji}\,y^{\prime}_{j}+b^{u}_{j}\\[5.69054pt] r_{i}=\sum_{j=1}^{m+n}a^{r}_{ji}\,y_{j}+b^{r}_{j}.\end{array}$ (62)
Here the vector y=[h,x]y\,{=}\,[h,x] is defined as before, and vector $y^{\prime}\,{=}\,[\sigma(r)\odot{}h,x]$. The RG occurring in $y^{\prime}$ determines how state $h$ is used in the update part $\tau(u_{i})$ of the ODE. The LRC’s time constant (TC) thus consists of a liquid-resistance liquid-capacitance, while the GRU’s TC consists of a liquid capacitance, only. As a consequence, GRUs have a less expressive TC but more expressive update.
0.A.3 Experiments for LRCs and LRCUs
In this section, we provide additional implementation details about our experimental evaluation of LRC in wide range of ODE modeling tasks and LRCUs on the popular time-series benchmarks for gated recurrent units. The benchmarks considered were the Localization Data for Person Activity, the IMDB Movie Sentiment Classification, the Permuted Sequential MNIST, and a relatively complex autonoumous-driving in the Lane-Keeping Task.
In general, LRC(U)s have more trainable parameters per cell. Therefore, we are training the other models with more cells to ensure a fair comparison between them. For the experiments, we determined the learning rate using the logarithmic scale of $\{10^{-4},10^{-3},10^{-2}\}$ by the validation loss. For the individual tests, see the more detailed setups below. They are run on Ubuntu 22.04 with Nvidia Tesla T4.
0.A.3.1 Neural ODE Experiments
For ODE modeling tasks, we used a 3-layer (units of 32, 32 and 2) Neural ODE and the LRC (units of 16) with additional input and output mapping, which hold 1k trainable parameters. The hyperparameters are presented in Table 7. We use sequences of 16 points during training and test the models on the whole sequences of 1000 data points giving them only the initial state information of $(x_{0},y_{0})$. As some ODE modeling tasks are more challenging, we allocated longer training iteration time for them. The training time is 0.4-0.5 seconds per iteration for both models.
@@ 0.A.3.1 Neural ODE Experiments
Iterations $1000/2000/4000$
  • Periodic Sinusoidal: $dx/dt=x\cdot(1-\sqrt{x^{2}+y^{2}})-y$, $dy/dt=x+y\cdot(1-\sqrt{x^{2}+y^{2}})$.
  • Spiral: $dx/dt=Ax$, where A=[[0.1,3],[3,0.1]A=[[-0.1,3],[-3,-0.1].
  • Duffing Oscillation: $dx/dt=y$, $dy/dt=x-x^{3}$.
  • Periodic Lotka-Volterra: $dx/dt=a\cdot x-b\cdot x\cdot y$, $dy/dt=-c\cdot y+d\cdot x\cdot y$ with $a=1.5$, $b=1$, $c=3$ and $d=1$.
  • Asymptotic Lotka-Volterra: $dx/dt=x\cdot(1-x)-x\cdot y$, $dy/dt=-y+d\cdot x\cdot y$ with $d=2$.
  • Non-linear Lotka-Volterra: $dx/dt=x\cdot(1-x)-a\cdot x\cdot y$, $dy/dt=y\cdot(1-y)+x\cdot y$ with $a=0.33$.
These examples originate from the TensorFlow implementation of [7], tfdiffeq.
0.A.3.2 Localization Data for Person Activity
@@
For LSTMs, GRUs, and MGUs we concatenate the time-step information (difference in time between to inputs) directly with the input features. This time step is used as the $\Delta_{t}$ value in LRCUs. LSTM, GRU and MGU networks contain 100 units, while LRCUs have 64 units.
@@
0.A.3.3 IMDB Movie Sentiment Classification
@@
Par. 40k 30k 20k 20k 20k
0.A.3.4 Permuted Sequential MNIST
@@
During the training of this task, the computed loss values of MGUs became NaNs in 2 out of the 3 runs in the middle of the experiments.
@@
0.A.3.5 Lane-Keeping Task
@@
Table 14: The shape and size of the layers in the convolutional head of the RNN-policy used to solve the Lane-Keeping Task. Settings are adapted from [12].
@@
Conv2D Filters: 24, Kernel size: 5, Stride: 2
Conv2D Filters: 36, Kernel size: 5, Stride: 1
@@
Conv2D Filters: 48, Kernel size: 3, Stride: 1
@@
Conv2D Filters: 64, Kernel size: 3, Stride: 1
@@
Conv2D Filters: 64, Kernel size: 3, Stride: 1
@@
Table 15: Hyperparameters of the Lane-Keeping experiment, adapted from [12].
@@
Optimizer AdamW
@@
0.A.4 Interpretability Experiments for LRCUs
In this section we provide additional details for the interpretability experiments for the Lane-Keeping Task.
Figure 10 shows our results in the summer season, where lighter-highlighted regions indicate the attention of the network. We found that the LSTM takes into account irrelevant parts of the image, during its decision-making. The rest of the models have most of the attention on the road, and LRCU networks especially, focus on the horizon. In Figure 11, we provide the attention maps for the winter season.
LSTM GRU MGU LRCU-A LRCU-S
0.0
0.1
0.2
@@ 0.A.4 Interpretability Experiments for LRCUs
By measuring the change of attention, we also compute the Structural Similarity Index (SSIM) [35] of the images pairwise, between the noise-free and the noisy saliency map for each model, respectively. This technique assesses the similarity between two images, with a focus on image degradation. An SSIM value of one denotes full similarity, whereas one of zero, denotes zero similarity. In our particular case, where we are comparing attention-map images, values close to one mean that the extra noise leads to less distortion in the attention map, which is a very desirable aim in the development of robust controllers. The SSIM value for 160016001600{-}1600 pairs of images, used in the attention-map comparisons between noise-free and noisy, summer and winter images, respectively, are presented in Figure 8. All results in this table are averaged over three runs for each model. As one can see, LRCUs demonstrate much more robust attention in the presence of Gaussian noise.
In Figure 12 we provide the neural activity of two cells of the lane-keeping policy for the winter season. As one can clearly see, the LRCU-S has a very gentle varying behavior closely matching the road geometry: Cell 1 is responsible for turning left, and Cell 2 is responsible for turning right.
@@
Refer to caption Refer to caption
@@
Refer to caption Refer to caption
Figure 12: Interpretability. The neural activity of two cells in the learned policy for the Lane-Keeping Task, projected over time on the 1km road driven in winter, for three popular gated recurrent units (LSTMs, GRUs, and MGUs) and for LRC units (LRCU-A, and LRCU-S). It is very hard to visually match the neural activity of LSTM, GRU, and MGU cells to the traversed road geometry, respectively. However, LRCU-S cells especially, display a very gentle varying and identifiable pattern during driving: Cell 1 is responsible for turning right, and Cell 2 for turning left.